Aspen Company estimates its manufacturing overhead to be $608,400 and its direct labor costs to be $507,000 for year 2. Aspen worked on three jobs for the year. Job 2-1, which was sold during year 2, had actual direct labor costs of $272,000. Job 2-2, which was completed, but not sold at the end of the year, had actual direct labor costs of $340,000. Job 2-3, which is still in work-in-process inventory, had actual direct labor costs of $68,000. Actual manufacturing overhead for year 2 was $802,700. Manufacturing overhead is applied on the basis of direct labor costs.

Required:

Prepare an entry to allocate over- or underapplied overhead to Work in Process, Finished Goods and Cost of Goods Sold. (If no entry is required for a transaction/event, select "No journal entry required" in the first account field.)

Journal entry worksheet

Transaction Index :

  • Record the allocation of over- or underapplied overhead.

Note: Enter debits before credits.

Transaction General Journal Debit Credit
1 Applied manufacturing overhead
Work-in-process inventory
Finished goods inventory
Cost of good sold

+
Answer #1
Predetermined Overhead rate 120% =608400/507000
Applied manufacturing overhead:
Job 2-1 326400 =272000*120%
Job 2-2 408000 =340000*120%
Job 2-3 81600 =68000*120%
Applied manufacturing overhead 816000
Over-applied manufacturing overhead 13300 =816000-802700
General Journal Debit Credit
Applied manufacturing overhead 816000
       Work­ in­ process inventory 5320 =13300/816000*326400
       Finished goods inventory 6650 =13300/816000*408000
       Cost of goods sold 1330 =13300/816000*81600
       Manufacturing overhead control 802700